OpenStudy (nightmarenight):

A right rectangular prism has a volume of 40 cm^3. Find the new volume of the prism if only the length of the prism is doubled.

OpenStudy (triciaal):

what is the formula for the volume of a rectangular prism?

OpenStudy (bee_see):

no picture? Volume of rectangular prism=(w)(l)(h). Here, 40=(w)(l)(h)

OpenStudy (triciaal):

let the length = l Base area =l* w double means 2 times if the length is doubled then the base area = (2l)* w

OpenStudy (nightmarenight):

@Bee_see no picture was attached to this question

OpenStudy (triciaal):

really don't need one

OpenStudy (nightmarenight):

(21)*w isn't an option o:

OpenStudy (triciaal):

important to see the formula and see the change

OpenStudy (triciaal):

that was not the answer!! that was for the new base area

OpenStudy (triciaal):

the volume for the prism = Bh where B is the base area and h is the height

OpenStudy (nightmarenight):

Ohhhh

OpenStudy (triciaal):

the area of the rectangle is l*w ; length * width the new area = (2l)* w the new volume = new base area * height = (2l)*w*h

OpenStudy (triciaal):

now compare the old volume with the new volume to see the effective change with the change in only the length

OpenStudy (nightmarenight):

Oh it's B

OpenStudy (triciaal):

I have no idea . I did not see the options. but do you see how to apply the formula?

OpenStudy (nightmarenight):

40cm^3

OpenStudy (triciaal):